# On the universality of the halo mass function beyond $\Lambda$CDM cosmology

### v1

@@ Abstract
The abundance of dark matter haloes as a function of halo mass is a key diagnostic for constraining the cosmological model. The theoretical framework based on excursion set arguments, when applied to an initial Gaussian random field of density fluctuations, predicts universal behaviour for this quantity, when variables are recast in terms of peak height. The great advantage of this, if true, is that it implies one simply needs to accurately simulate only a single cosmological model to build an emulator for any other cosmology of interest. This tantalising possibility has inspired a number of studies over the years. In practice, the diversity of ways for defining haloes has led to a variety of mixed results concerning this issue. In this work, we utilise a suite of high-resolution cosmological $N$-body simulations, to revisit this question for friends-of-friends haloes. We perform our study in the context of the flat, time evolving dark energy model (hereafter $w$CDM), and with simple modifications of the primordial physics afforded through variations of the scalar power spectral index and its possible running. We construct the universal mass function locus from our fiducial simulation (a $\Lambda$CDM model) and emulate this using a linear interpolating function. We then compare this against the loci that we measure for our array of alternate models. We find mass functions that are consistent with universality to within $\lesssim 5\%$ in the fractional difference, with respect to variations of the 8 cosmological parameters that we have considered (2 variations per parameter) and for redshifts $z<7$.
@@
pubyear: 2024pagerange: On the universality of the halo mass function beyond $\Lambda$CDM cosmologyC
@@ 1 Introduction
One of the keystones of modern cosmology is that the majority of gravitating mass in the Universe, some $\sim 85\%$, is in the form of collisionless cold dark matter (hereafter CDM) with the remainder being in the form of the known standard model particles ‘baryons’ and neutrinos (Planck Collaboration et al., 2014). The true physical nature of the dark matter particle remains a mystery (for a review see Bertone, 2010). Primordial fluctuations in the CDM are thought to be sourced after a period of inflation, whereupon they are well described by a Gaussian random field with a power spectrum that has almost scale-invariant potential fluctuations. The wealth of late-time cosmic structures that we observe today, the sheets, filaments, and clusters – the cosmic web – then forms subsequently through the gravitational instability of these primordial density seeds.
@@
One of the commonly used theoretical frameworks to describe the late-time distribution of matter is the halo model. This simply divides all of the matter into a set of distinct units – haloes (see, Cooray & Sheth, 2002; Asgari et al., 2023, for reviews). If one considers the halo model in its basest form, where all particles are considered to belong to a halo, going down to single particle haloes, then it is simply an equivalence class on the matter. However, to elevate this from a tautology and so make useful predictions one requires a model for the abundance of haloes, their evolution with time, their clustering, and their internal density distributions. With these in hand one can make predictions for an astonishingly wide array of cosmological observables (for examples of applications see Cooray & Sheth, 2002; Asgari et al., 2023). The halo model framework has been of great utility for many of the wide field large-scale structure surveys (especially if one encompasses SHAMs and HAMs etc). For example, recently in the Dark Energy Survey it was extensively used in Friedrich et al. (2021) for building covariance models, and for the KiDS survey by Dvornik et al. (2023) who explored the galaxy matter connection, among others. It will no doubt also play an important role in the exploitation of the upcoming surveys like DESI (DESI Collaboration et al., 2016), Euclid (Laureijs et al., 2011) and LSST (Ivezić et al., 2019) surveys. Thus to fully exploit the data from these missions, there is high demand for theoretical models that can accurately predict the abundance of dark matter haloes over a wide range of masses and cosmic time scales to high precision. And doing so for a wide array of viable scenarios beyond the vanilla flavoured vacuum energy dominated CDM model (here after $\Lambda$CDM). Exploring this is one of the aims of this paper.
@@
An analytic form of the halo mass function (HMF) was first derived by Press & Schechter (1974) by means of applying the spherical collapse model to arbitrary points in the initial linear matter density field. They noted that their model predicted that only half of the mass in the Universe was in collapsed haloes, which they corrected in an ad-hoc fashion by multiplication of their final result by 2! Later on, Bond et al. (1991) provided a robust derivation of their formula, which included the missing factor, using excursion set theory. Interestingly, it was found that the Press-Schechter HMF could be rearranged into a universal form, i.e., being completely independent of cosmology and redshift, through a judicious choice of variables. Although the Press-Schechter function failed to accurately predict the number densities of very large and small mass haloes that were measured in cosmological simulations (Lacey & Cole, 1994; Sheth & Tormen, 1999; Jenkins et al., 2001), the possibility of a universal HMF inspired a number of subsequent studies.
@@
  • FoF Haloes: One of the earliest operational definitions was given through the Friends-of-Friends (FoF) algorithm of Davis et al. (1985). This identifies haloes in numerical simulations as all linked particles with separations less than $b\bar{n}^{-1/3}$, where $\bar{n}$ is the mean particle density and $b$ is the linking parameter typically set to $b=0.2$. This halo finder makes no assumptions about the shape of the halo and is computationally efficient, but can suffer from ‘bridging’ between neighbouring groups. Subsequent studies of FoF haloes show that the mass functions of these objects can exhibit, approximate, universal behaviour (Sheth & Tormen, 1999; Jenkins et al., 2001; White, 2002; Reed et al., 2003; Warren et al., 2006; Reed et al., 2007; Lukić et al., 2007; Crocce et al., 2010; Manera et al., 2010; Courtin et al., 2011; Bhattacharya et al., 2011; Juan et al., 2014).
  • SO haloes: Another common halo finder, that is possibly more physically motivated, is the spherical overdensity (SO) method, which identifies haloes as all matter within a sphere whose average density is some fraction $\Delta$ of the background density (Lacey & Cole, 1994). This was inspired by the spherical collapse model, for which, in the Einstein de-Sitter model, after virialisation the average overdensity is found to be $\sim 180\rho_{\rm crit}$, where $\rho_{\rm crit}$ is the critical density. In many studies the characteristic overdensity is typically fixed at $\Delta=200$ for all cosmological model, this is in contrast to the spherical collapse density which varies with the underlying model (Lacey & Cole, 1993; Eke et al., 1996). Subsequent numerical studies of SO haloes show that the mass functions of these haloes departs from universal behaviour (Tinker et al., 2008; Watson et al., 2013; Despali et al., 2016; McClintock et al., 2019; Ondaro-Mallea et al., 2021).
  • Splashback haloes: More recently, another, perhaps even more physically-motivated, model for halo finding involves using the so called splashback radius (More et al., 2015). The splashback radius is located at the place where the accreted matter reaches its first orbital apocentre after turnaround, which corresponds to a sharp drop in the halo density profile. All matter within this radius is then considered to be part of the halo. A subsequent study, has shown that haloes defined using the splashback-radius method are also potentially good candidates in favour of universality (Diemer, 2020).
@@
Table 1: Cosmological parameters used the Dämmerung simulations. Column 1 denotes the simulation series. The next 8 columns denote the 8 cosmological parameters listed in Eq. (1) that define each simulated model: columns (2) and (3) give $w_{0}$ and $w_{a}$ which are parameters used to define the dark energy equation of state; column (4) gives the density parameter for dark energy $\Omega_{\rm DE}$; columns (5) and (6) give the physical densities of CDM $\omega_{\rm c}=\Omega_{\rm c}h^{2}$ and baryons $\omega_{\rm b}=\Omega_{\rm b}h^{2}$; columns (7), (8) and (9) give the primordial amplitude of scalar perturbations $A_{\rm s}$, the spectral index $n_{\rm s}$ and the running of the spectral index $\alpha_{\rm s}$. Columns (10)-(12) give $\Omega_{\rm m}$, $h$ and $\sigma_{8}$, all of which are derived parameters. Parameter values that are deviations from the fiducial model are highlighted in bold.
Para $w_{0}$ $w_{a}$ $\Omega_{\rm DE}$ $\omega_{\rm c}$ $\omega_{\rm b}$ $A_{\rm s}[\times 10^{-9}]$ $n_{\rm s}$ $\alpha_{\rm s}$ $\Omega_{\rm m}$ $h$ $\sigma_{8}$
@@ 2 Numerical simulations
Hence, the matter power spectrum can be written in terms of primordial quantities as:
@@
Note that for this parameterisation, taken along with the assumption of flatness, we can derive a number of other useful cosmological parameters: the matter density parameter $\Omega_{\rm m}=1-\Omega_{\rm DE}$; the dimensionless Hubble rate $h=\sqrt{\left[\omega_{\rm b}+\omega_{\rm c}\right]/\Omega_{\rm m}}$; the square-root of the variance of matter fluctuations in spheres of radius $R=8\ h^{-1}{\rm Mpc}$ can also be computed $\sigma_{8}$ – for details of this see Eq. (15). Through varying each of the 8 cosmological parameters in Eq. (1) positively and negatively, additional 16 variational models are included to investigate the dependence of the nonlinear structure formation on these parameters. Note that these variations were chosen to facilitate the construction of derivatives of the observables with respect to the cosmological parameters and so enable Taylor expansions of the model around the fiducial point to be constructed. The exact values of the cosmological parameters used in all models are listed in Table 1. Note that here we correct typos for the parameters that were presented in Table 1 of Smith & Angulo (2019) for models listed as V5 and V6, which varied the dark energy density (or matter density).
@@
All simulation runs were performed using a special version of the parallel Tree-PM code Gadget-3, which was developed for the purpose of generating the Millennium-XXL simulation (Springel, 2005; Angulo et al., 2012). Each run was performed with $N_{\rm p}=2048^{3}$ dark matter particles and $N_{\rm grid}=2048^{3}$ PM grid in a box of comoving size $L_{\rm box}=500\ h^{-1}\rm{Mpc}$. The particle mass for all simulations was $m_{\rm p}=1.240\times 10^{9}\ h^{-1}M_{\odot}$, with the exception of models V5 and V6 (see Table 1), which have $m_{\rm p}=1.038\times 10^{9}\ h^{-1}M_{\odot}$ and $1.442\times 10^{9}\ h^{-1}M_{\odot}$, respectively, arising due to the variation of $\Omega_{\rm m}$. Note that we here correct also the typos of the particle masses listed in Table 2 of Smith & Angulo (2019). The softening length was set to $l_{\rm soft}=7.5h^{-1}\rm kpc$, which is about $3\%$ of the mean inter-particle spacing $l=L_{\rm box}/\sqrt[3]{N_{\rm p}}$, which is widely used as the empirical optimal softening length (Power et al., 2003).
@@
The fiducial model consists of 10 realisations with the same initial power spectrum but different realisations of the Gaussian random field. These were generated using a modified version of the public 2LPT algorithm of (Crocce et al., 2006). By contrast, only one run was performed for each of the variational models, and the initial Gaussian random field generation was set for each variation to match that of run 1 in the fiducial model with the exception of the power spectrum model and linear growth rates which were modified accordingly for each model. This was done to help minimise the effect of large-scale cosmic variance when making comparisons between variations. Note that we refer to run 1 as the fiducial run hereafter unless otherwise stated. The haloes were identified using the on-the-fly FoF algorithm built into Gadget-3 and was run with the linking length parameter set to $b=0.2$. 63 snapshots were output between $z=49$ and $z=0$ with a hybrid linear-logarithmic time spacing. In this study we primarily focus on 8 snapshots selected at $z=0.00$, $0.51$, $0.99$, $1.50$, $2.05$, $3.08$, $4.54$, and $6.75$.
@@ 3.1 Background theory
$\nu f(\nu)=\frac{m}{\overline{\rho}_{0}}[mn(m,z)]\frac{{\rm d}\log{m}}{{\rm d}\log{\nu}},\ .$ (8)
@@
Lacey & Cole (1994) and later Sheth & Tormen (1999) subsequently found that the Press-Schechter model overestimated the abundance of low-mass haloes and underrepresented the abundance of high-mass haloes when compared to numerical simulations. This inspired Sheth & Tormen (1999) and Sheth et al. (2001) to explore ellipsoidal collapse and more complicated excursion set scenarios. This opened the path to more complex possibilities for the analytic form of the HMF. However, when cast in terms of the multiplicity function, they retained the universal structure, e.g. Sheth & Tormen (1999) proposed:
@@
where $a$ and $p$ were related to the ellipsoidal collapse scenario, but were taken as free parameters, and $A$ can be determined by requiring $\int f(\nu){\rm d}\nu=1$. This function was shown to give a significantly better analytic approximation to simulated data. This approach paved the way for a wide array of numerical studies of the HMF (Jenkins et al., 2001; White, 2002; Reed et al., 2003; Warren et al., 2006; Reed et al., 2007; Lukić et al., 2007; Tinker et al., 2008; Crocce et al., 2010; Bhattacharya et al., 2011; Courtin et al., 2011; Reed et al., 2013; Watson et al., 2013; Castorina et al., 2014; Bocquet et al., 2016; McClintock et al., 2019; Euclid Collaboration et al., 2023)  [^3] Note that we view studies that have chosen to model the mass function using a function $f(\sigma^{-1})$ la Jenkins et al. (2001) as being approximately equivalent to those modelling $\nu$, since $\delta_{\rm c}$ is a very weak function of cosmology and since they encode the same information, but see Courtin et al. (2011) for further discussion of this issue.

### v2

@@ Abstract
Theoretical frameworks based on Press-Schechter formalism and excursion set arguments suggest that the abundance of dark matter haloes exhibits universal behaviour when expressed in terms of peak height. If true, this implies that a single high-accuracy cosmological simulation could serve as a basis for constructing an emulator applicable to any other cosmology of interest. This tantalising possibility has inspired numerous studies over the years. However, in practice, different ways of defining haloes have led to mixed results concerning this issue. In this work, we utilise a suite of high-resolution cosmological $N$-body simulations, to revisit this question for friends-of-friends haloes under the flat time-evolving $w$CDM model, with simple modifications of the primordial physics via variations in the scalar spectral index and its running. We construct a reference locus of $\nu f(\nu)$ from our fiducial $\Lambda$CDM simulation and compare it against measurements from alternative models. We find that deviations from the locus remain within $5\%$ when varying each of the parameters within the ranges: ${w_{0}}=-1.0\pm 0.1$, $w_{a}=0\pm 0.2$, $\Omega_{\rm DE}=0.693\pm 0.050$, $\omega_{\rm c}=0.119\pm 0.006$, $\omega_{\rm b}=0.0222\pm 0.0011$, $A_{\rm s}=(2.15\pm 0.22)\times{10^{-9}}$, $n_{\rm s}=0.961\pm 0.048$, $\alpha_{\rm s}\ =0\pm 0.01$, for redshift $z<7$.
@@
pubyear: 2025pagerange: On the universality of the halo mass function beyond $\Lambda$CDM cosmology14
@@ 1 Introduction
One of the keystones of modern cosmology is that the majority of gravitating mass in the Universe, some $\sim 85\%$, is in the form of collisionless cold dark matter (hereafter CDM) with the remainder being in the form of the known standard model particles ‘baryons’ and neutrinos (Planck Collaboration et al., 2014). The true physical nature of the dark matter particle remains a mystery (for a review see Bertone, 2010). Primordial fluctuations in the CDM are thought to be sourced after a period of inflation, whereupon they are well described by a Gaussian random field with a power spectrum that has almost scale-invariant potential fluctuations. However, during the radiation-dominated era, fluctuations inside the horizon are suppressed due to the presence of radiation pressure, shaping the matter power spectrum we observe today. This suppression results in a characteristic turnover in the power spectrum, with a spectral slope approaching -3 on small scales rather than 1. The wealth of late-time cosmic structures that we observe today, the sheets, filaments, and clusters – the cosmic web – then forms subsequently through the gravitational instability of these processed primordial density seeds.
@@
One of the commonly used theoretical frameworks to describe the late-time distribution of matter is the halo model. This simply divides all of the matter into a set of distinct units – haloes (see, Cooray & Sheth, 2002; Asgari et al., 2023, for reviews). If one considers the halo model in its basic form, where all particles are considered to belong to a halo, going down to single particle haloes, then it is simply an equivalence class on the matter. However, to elevate this from a tautology and so make useful predictions one requires a model for the abundance of haloes, their evolution with time, their clustering, and their internal density distributions. With these in hand one can make predictions for an astonishingly wide array of cosmological observables (for examples of applications see Cooray & Sheth, 2002; Asgari et al., 2023). The halo model framework has been of great utility for many of the wide field large-scale structure surveys (especially if one encompasses SHAMs and HAMs etc). For example, recently in the Dark Energy Survey it was extensively used in Friedrich et al. (2021) for building covariance models, and for the KiDS survey by Dvornik et al. (2023) who explored the galaxy matter connection, among others. It will no doubt also play an important role in the exploitation of the upcoming surveys like DESI (DESI Collaboration et al., 2016), Euclid (Laureijs et al., 2011) and LSST (Ivezić et al., 2019) surveys. Thus to fully exploit the data from these missions, there is high demand for theoretical models that can accurately predict the abundance of dark matter haloes over a wide range of masses and cosmic time scales to high precision. And doing so for a wide array of viable scenarios beyond the vanilla flavoured vacuum energy dominated CDM model (hereafter $\Lambda$CDM). Exploring this is one of the aims of this paper.
@@
An analytic form of the halo mass function (HMF) was first derived by Press & Schechter (1974) by means of applying the spherical collapse model to arbitrary points in the initial linear matter density field. They noted that their model predicted that only half of the mass in the Universe was in collapsed haloes, which they corrected in an ad hoc fashion by multiplication of their final result by 2! Later on, Bond et al. (1991) provided a robust derivation of their formula, which included the missing factor, using excursion set theory. Interestingly, it was found that the Press-Schechter HMF could be rearranged into a universal form, i.e., being completely independent of cosmology and redshift, through a judicious choice of variables. Although the Press-Schechter function failed to accurately predict the number densities of very large and small mass haloes that were measured in cosmological simulations (Lacey & Cole, 1994; Sheth & Tormen, 1999; Jenkins et al., 2001), the possibility of a universal HMF inspired a number of subsequent studies.
@@
  • FoF haloes: One of the earliest operational definitions was given through the Friends-of-Friends (FoF) algorithm of Davis et al. (1985). This identifies haloes in numerical simulations as all linked particles with separations less than $b\bar{n}^{-1/3}$, where $\bar{n}$ is the mean particle density and $b$ is the linking parameter typically set to $b=0.2$. This halo finder makes no assumptions about the shape of the halo and is computationally efficient and fast. The method directly groups particles without needing to smooth densities, estimate continuous fields, or iteratively determine the centre and boundary of the halo, but can suffer from ‘bridging’ between neighbouring groups. Subsequent studies of FoF haloes show that the mass functions of these objects can exhibit, approximate, universal behaviour (Sheth & Tormen, 1999; Jenkins et al., 2001; White, 2002; Reed et al., 2003; Warren et al., 2006; Reed et al., 2007; Lukić et al., 2007; Crocce et al., 2010; Manera et al., 2010; Courtin et al., 2011; Bhattacharya et al., 2011; Juan et al., 2014).
  • SO haloes: Another common halo finder, that is possibly more physically motivated, is the spherical overdensity (SO) method, which identifies haloes as all matter within a sphere whose average density is some fraction $\Delta$ of the background density (Lacey & Cole, 1994). This was inspired by the spherical collapse model, for which, in the Einstein de-Sitter model, after virialisation the average overdensity is found to be $\sim 180\rho_{\rm crit}$, where $\rho_{\rm crit}$ is the critical density. In many studies the characteristic overdensity is typically fixed at $\Delta=200$ for all cosmological model, this is in contrast to the spherical collapse density which varies with the underlying model (Lacey & Cole, 1993; Eke et al., 1996). On the other hand, in practice the algorithms using SO approach to define halo boundaries may differ in identifying the centre of halo (see, Euclid Collaboration et al., 2023, for a review). Subsequent numerical studies of SO haloes show that with the exception of virial overdensity, the mass functions of these haloes more likely departs from universal behaviour (Tinker et al., 2008; Watson et al., 2013; Despali et al., 2016; McClintock et al., 2019; Ondaro-Mallea et al., 2021; Gavas et al., 2023) in Einstein-de Sitter and $\Lambda$CDM cosmologies. In contrast, a different trend was observed in modified gravity scenarios (Gupta et al., 2022).
  • Other haloes: More recently, another two, perhaps even more physically-motivated, models for halo finding were proposed. One involves using the caustic splashback radius (Diemer & Kravtsov, 2014; Adhikari et al., 2014; More et al., 2015). The splashback radius is located at the place where the accreted matter reaches its first orbital apocentre after turnaround, which corresponds to a sharp drop in the halo density profile – in the case of perfect spherical symmetry forming a zero density caustic. All matter within this radius is then considered to be part of the halo. The other one defines halo as a collection of particles orbiting in their own self-generated potential (Garcia et al., 2023). A subsequent study, has shown that haloes defined using the splashback-radius method are also potentially good candidates in favour of universality (Diemer, 2020).
@@
Table 1: Cosmological parameters used in the Dämmerung simulations. Column 1 denotes the simulation series. The next 8 columns denote the 8 cosmological parameters listed in Eq. (1) that define each simulated model: columns (2) and (3) give $w_{0}$ and $w_{a}$ which are parameters used to define the dark energy equation of state; column (4) gives the density parameter for dark energy $\Omega_{\rm DE}$; columns (5) and (6) give the physical densities of CDM $\omega_{\rm c}=\Omega_{\rm c}h^{2}$ and baryons $\omega_{\rm b}=\Omega_{\rm b}h^{2}$; columns (7), (8) and (9) give the primordial amplitude of scalar perturbations $A_{\rm s}$, the spectral index $n_{\rm s}$ and the running of the spectral index $\alpha_{\rm s}$. Columns (10)-(12) give $\Omega_{\rm m}$, $h$ and $\sigma_{8}$, all of which are derived parameters. Parameter values that are deviations from the fiducial model are highlighted in bold.
Parameter $w_{0}$ $w_{a}$ $\Omega_{\rm DE}$ $\omega_{\rm c}$ $\omega_{\rm b}$ $A_{\rm s}[\times 10^{-9}]$ $n_{\rm s}$ $\alpha_{\rm s}$ $\Omega_{\rm m}$ $h$ $\sigma_{8}$
@@ 2 Numerical simulations
Hence, the matter power spectrum can be written in terms of primordial quantities as (Smith & Simon, 2025):
@@
Note that for this parameterisation, taken along with the assumption of flatness, we can derive a number of other useful cosmological parameters: the matter density parameter $\Omega_{\rm m}=1-\Omega_{\rm DE}$; the dimensionless Hubble rate $h=\sqrt{\left[\omega_{\rm b}+\omega_{\rm c}\right]/\Omega_{\rm m}}$; the square-root of the variance of matter fluctuations in spheres of radius $R=8\ h^{-1}{\rm Mpc}$ can also be computed $\sigma_{8}$ – for details of this see Eq. (15), and varying $A_{\rm s}$ is equivalent to varying $\sigma_{8}^{2}$. Through varying each of the 8 cosmological parameters in Eq. (1) positively and negatively, additional 16 variational models are included to investigate the dependence of the nonlinear structure formation on these parameters. Note that these variations were chosen to facilitate the construction of derivatives of the observables with respect to the cosmological parameters and so enable Taylor expansions of the model around the fiducial point to be constructed. The exact values of the cosmological parameters used in all models are listed in Table 1. Note that here we correct typos for the parameters that were presented in Table 1 of Smith & Angulo (2019) for models listed as V5 and V6, which varied the dark energy density (or matter density).
Figure 1: Fractional difference in the linear theory power spectra at $z=0$ between the variational models and the fiducial model. The left panel shows the variations in $w_{0},w_{a},\Omega_{\rm DE},\omega_{\rm c}$, and the right panel $\omega_{\rm b},A_{\rm s},n_{\rm s},\alpha_{\rm s}$. Solid coloured lines represent positive parameter variations, and dashed lines indicate negative variations.
@@
The linear theory matter power spectra used to generate the initial conditions (ICs) were generated at $z=0$ for all models using the Einstein-Boltzmann solver code CAMB. These were then rescaled back to $z=49$ as the input linear power spectra using the linear growth factor described in Section 3.3 in Smith & Angulo (2019), to feed a modified version of the public 2LPT algorithm of (Crocce et al., 2006). Figure 1 shows the fractional difference in the linear power spectra at $z=0$ between the variational models and fiducial model. All simulation runs were performed using a special version of the parallel Tree-PM code Gadget-3, which was developed for the purpose of generating the Millennium-XXL simulation (Springel, 2005; Angulo et al., 2012). Each run was performed with $N_{\rm p}=2048^{3}$ dark matter particles and $N_{\rm grid}=2048^{3}$ PM grid in a box of comoving size $L_{\rm box}=500\ h^{-1}\rm{Mpc}$. The particle mass for all simulations was $m_{\rm p}=1.240\times 10^{9}\ h^{-1}M_{\odot}$, with the exception of models V5 and V6 (see Table 1), which have $m_{\rm p}=1.038\times 10^{9}\ h^{-1}M_{\odot}$ and $1.442\times 10^{9}\ h^{-1}M_{\odot}$, respectively, arising due to the variation of $\Omega_{\rm m}$. Note that we here correct also the typos of the particle masses listed in Table 2 of Smith & Angulo (2019). The softening length was set to $l_{\rm soft}=7.5h^{-1}\rm kpc$, which is about $3\%$ of the mean inter-particle spacing $l=L_{\rm box}/\sqrt[3]{N_{\rm p}}$, which is widely used as the empirical optimal softening length (Power et al., 2003).
@@
The fiducial model consists of 10 realisations with the same initial power spectrum, but different realisations of the Gaussian random field. We note that the fields are true Gaussian random fields, with the initial Fourier modes having random phases and the amplitudes being drawn from the Rayleigh distribution. For the variational models, only one run was performed, and for each variation the initial Gaussian random field generation was set for each variation to match that of run 1 in the fiducial model with the exception of the power spectrum model and linear growth rates, which were modified accordingly for each model. This was done to help minimise the effect of large-scale cosmic variance when making comparisons between variations. Note that we refer to run 1 as the fiducial run hereafter unless otherwise stated. The haloes were identified using the on-the-fly FoF algorithm built into Gadget-3 and was run with the linking length parameter set to $b=0.2$. 63 snapshots were output between $z=49$ and $z=0$ with a hybrid linear-logarithmic time spacing. In this study we primarily focus on 8 snapshots selected at $z=0.00$, $0.51$, $0.99$, $1.50$, $2.05$, $3.08$, $4.54$, and $6.75$.
@@ 3.1 Background theory
$\nu f(\nu)=\frac{m}{\overline{\rho}_{0}}[mn(m,z)]\frac{{\rm d}\log{m}}{{\rm d}\log{\nu}}\ .$ (8)
@@
Lacey & Cole (1994) and later Sheth & Tormen (1999) subsequently found that the Press-Schechter model overestimated the abundance of low-mass haloes and underrepresented the abundance of high-mass haloes when compared to numerical simulations. This inspired Sheth et al. (2001) to explore ellipsoidal collapse and more complicated excursion set scenarios. This opened the path to more complex possibilities for the analytic form of the HMF. However, when cast in terms of the multiplicity function, they retained the universal structure, e.g. Sheth & Tormen (2002) proposed:
@@
where $a$ and $p$ were related to the ellipsoidal collapse scenario, but were taken as free parameters, and $A$ can be determined by requiring $\int f(\nu){\rm d}\nu=1$. This function was shown to give a significantly better analytic approximation to simulated data. This approach paved the way for a wide array of numerical studies of the HMF (Jenkins et al., 2001; White, 2002; Reed et al., 2003; Warren et al., 2006; Reed et al., 2007; Lukić et al., 2007; Tinker et al., 2008; Crocce et al., 2010; Bhattacharya et al., 2011; Courtin et al., 2011; Reed et al., 2013; Watson et al., 2013; Castorina et al., 2014; Bocquet et al., 2016; Del Popolo et al., 2017; McClintock et al., 2019; Bocquet et al., 2020; Euclid Collaboration et al., 2023)  [^3] Note that we view studies that have chosen to model the mass function using a function $f(\sigma^{-1})$ la Jenkins et al. (2001) as being approximately equivalent to those modelling $\nu$, since $\delta_{\rm c}$ is a very weak function of cosmology and since they encode the same information, but see Courtin et al. (2011) for further discussion of this issue.